OpenStudy (xmoses1):

y varies inversely as x, and y = 50 when x = 10. What is the value of y when x = 20? 100 25 10

OpenStudy (the_fizicx99):

y = k/x 50 = k/10 *10 *10 k = 500 y = 500/20 y = 25

OpenStudy (xmoses1):

oh thanks i though the answer was 100 O.o thanks for helping me out!